OpenStudy (anonymous):

What is the equation of the line in slope-intercept form that passes through the point (–2, 4) and is perpendicular to the line y = –2x + 4? Answer y = negative one-halfx – 5 y = negative one-halfx – 2 y = one-halfx + 5 y = one-halfx + 2

OpenStudy (anonymous):

y = one-halfx + 5 (–2, 4) and is perpendicular to the line y = –2x + 4? perpendicular slope m = 1/2 y - 4 = (1/2 ) ( x + 2) y = x/2 + 1 + 4 = x/2 + 5